OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x^4 + 4x^2

OpenStudy (anonymous):

is the answer f(x) = x^2 (x + 2i)(x - 2i) ?

OpenStudy (amoodarya):

x^2(x^2+4) =x^2(x+2i)(x-2i) yes